OpenStudy (anonymous):

Complete the first four steps for graphing the quadratic function y = -x^2 - 4x - 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you familiar with completing the square?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

ok I'll show you how to find the vertex

jimthompson5910 (jim_thompson5910):

y = -x^2 - 4x - 3 is in the form y = ax^2 + bx + c where a = -1 b = -4 c = -3

jimthompson5910 (jim_thompson5910):

we take a = -1, b = -4 and plug them into x = -b/(2a) to get x = -b/(2a) x = -(-4)/(2(-1)) x = 4/(-2) x = -2

jimthompson5910 (jim_thompson5910):

So the axis of symmetry is x = -2 That means the x coordinate of the vertex is -2

jimthompson5910 (jim_thompson5910):

Plug x = -2 back into the original equation to find the y coordinate of the vertex

jimthompson5910 (jim_thompson5910):

y = -x^2 - 4x - 3 y = -(-2)^2 - 4(-2) - 3 .. replace every x with -2 y = -4 - 4(-2) - 3 y = -4 + 8 - 3 y = 1

jimthompson5910 (jim_thompson5910):

The x coordinate of the vertex is x = -2, the y coordinate is y = 1 With me so far?